Kinkead Inc. forecasts that its free cash flow in the coming year, i.e., at t = 1, will be −$10 million, but its FCF at t = 2 will be $20 million. After Year 2, FCF is expected to grow at a constant rate of 4% forever. If the weighted average cost of capital is 14%, what is the firm's value of operations, in millions?


1
Expert's answer
2021-07-20T16:02:25-0400

EXPLANATION;- Free cash flow FCF 1 = -$10 million

t = 1

Free cash flow FCF 2= $20 million

t = 2

FCF grow rate = 4%

average cost of capital = 14%

Now to find out

what is the firm's value of operations,

SOLUTION -

first we get here firm value in year 2 that is express as

firm value in year 2 = expected FCF in 3 ÷ (cost of capital - growth)  .........1

put here value

firm value in year 2 = "\frac{20*(1+0.04)}{0.14-0.04}"

firm value in year 2 = 208 million

and

firm value of operation this year will be as

firm value = discounted value in year 2 + discounted FCF1 and FCF2  .............2

firm value = "\frac{208}{(1+0.14)^2}+\frac{20}{(1+0.14)^2}+\frac{-10}{(1+0.14)}"

firm value = 166.67 = 167 million